OpenStudy (anonymous):

Please Help! Anyone who helps I will somehow be in debt with! I really need assistance here! But, I cannot participate in the discussion as much as I'd like. Please just try and show in at least some steps how to plot it, and the range.

jimthompson5910 (jim_thompson5910):

To graph this, you can plot a series of points and draw a curve through them or you can shift the parent graph y = 1/x one unit to the left

jimthompson5910 (jim_thompson5910):

You have the correct domain and the correct explanation for it

jimthompson5910 (jim_thompson5910):

Shifting y = 1/x to the left one unit is much easier

jimthompson5910 (jim_thompson5910):

To find the range, swap x and y to get x = 1/(y+1) Then solve for y x(y+1) = 1 xy+x = 1 xy = 1-x y = (1-x)/x So what value(s) of x aren't allowed here?

OpenStudy (anonymous):